jmoorene:

Find S20 for the series -1 + -3 + -5 + -7 +…

Ultrilliam:

@Vocaloid

sillybilly123:

I have a feeling that \(S_{20}\) means the series [ie SUM] of the first 20 numbers in the sequence. and not the 20th term of the sequence. For the **given 4 terms**: \(S_4 = -1 + -3 + -5 + -7\) We can also say, reversing the order, that: \(S_4 = -7 + -5 + -3 + -1\) And then add them: \(2 \times S_4 = -7 - 1 + -5 - 3 + -3 - 5 + -1 - 7 = -8 + -8 + - 8 + -8 = 4( - 8)\) it's -16 all the way but they want \(S_{20}\) so you work that out to \(S_{20}\), or generalise to \(S_n\) ...or you go look up a formula on Wiki for **arithmetic series**. either gets you the answer :))